Five hundred children participated in a field demonstration. Their heights averaged is 110 cm with a standard deviation of 6 cm. What is the probability that a child picked at random has a height less than 104 cm?


1
Expert's answer
2021-05-03T04:44:15-0400

We assume normal distribution with the mean ("\mu" ) and standard deviation ("\sigma" ) which are given as follows



"\mu =110 \text{ cm and } \sigma =6\text{ cm }."


Therefore,



"P (x < 104 ) = P (z<\frac{x-\bar{\mu}}{\sigma})=P(z<\frac{104-110}{6}),\\
=P(z<\frac{-6}{6})=P(z<-1)=0.15866."


Thus, the probability that a child picked at random has a height less than 104 cm is 0.15866.